To determine the Ca^2+ concentration in water samp
Answer #1

3 )

Ca+2 + EDTA ----------------> CaEDTA

millimoles of EDTA = millimoles of Ca+2

0.01988 x 15.80 = C x 25

C = 0.01256 M

molar concentration of Ca+2 = 0.01256 M

molarity = mass (g) / molar mass ) x 1/ V (L)

0.01256 mole/ litre = 0.01256 x 40.08 x g / L

= 0.5036 g / L

= 503.6 mg / L

= 503.6 ppm

in ppm Ca+2 concentration = 503.6 ppm

4 )

1) before titration starts

moles of benzoic acid = 0.2915 / 122.12 = 2.39 x 10^-3

molarity = moles / volume = 2.39 x 10^-3 / 100 x 10^-3 = 0.0239 M

C = 0.0239 , Ka = 6.3 x 10^-5

[H+] = sqrt (Ka x C) = sqrt (6.3 x 10^-5 x 0.0239) = 1.227 x 10^-3 M

pH = -log [H+] = -log (1.227 x 10^-3)

pH = 2.91

2) pH at half-equivalence point:

at this point . pH = pKa

pKa = -log Ka = -log (6.3 x 10^-5) = 4.20

pH = 4.20